significent @ Use the given values of n and p to find the minimum related binomial distribution. Round your answer to the nearest hundredth. I value μ-2σ and the maximum value μ + 2σ for the v x35 23) n-94, p 0.24 rignificant 23)
Answer #1

n = 94 and p = 0.24

mean = np = 94*0.24 = 22.5600
std. dev. = sqrt(94*0.24*0.76) = 4.1407

minimum value = mu - 2*sigma = 22.56 - 2*4.1407 = 14.2786
maximum value = mu + 2*sigma = 22.56 + 2*4.1407 = 30.8414

b)

35 is not part of the above CI, hence 35 is significant